OpenStudy (anonymous):

In a box there are 6 red, 5 green and 4 white balls. If four balls are selected at random and the color noted, find the probability that none are green given that none are white.

OpenStudy (anonymous):

i know the answer if that helps, i just need to know how to do it

OpenStudy (paxpolaris):

given that none are white ... so forget all the white balls. there are 6 red, 5 green balls. 4 balls are selected at random... Find the probability that none are green. ... that means find the probability that all are Red. .

OpenStudy (anonymous):

6/11?

OpenStudy (paxpolaris):

6/11 i the probability of the first ball you pick what's the probability for the next ball

OpenStudy (anonymous):

does it have to be red?

OpenStudy (paxpolaris):

yes' all 4 have to be red since none are green

OpenStudy (anonymous):

okay, so i THINK ... 5/10

OpenStudy (paxpolaris):

right ...

OpenStudy (anonymous):

if right then 1/2

OpenStudy (paxpolaris):

then 4/9 for the next and 3/8 for the fourth ball

OpenStudy (anonymous):

do you add or multiply?

OpenStudy (paxpolaris):

So you answer should be : \[\Large {6 \over 11} \times{5 \over 10} \times{4\over9} \times{3\over8}\]

OpenStudy (anonymous):

1/22 thanks :)